# Constraints on the $\gamma$-parameter for the vacuum solution of Cotton gravity with geodesics and shadows

### v1

@@ Abstract
We consider a recently introduced extension of General Relativity dubbed as Cotton gravity (CG), based on the use of the Cotton tensor, to estimate the size of a new constant $\gamma$ appearing within a spherically symmetric, vacuum solution of the theory. Taking into account its non-asymptotically flat character, we use the inferred size of the central brightness depression of the supermassive object at the heart of the Milky Way galaxy (Sgr A*) by the Event Horizon Telescope to constrain at $2\sigma$ the CG parameter as $\gamma M\approx 3.5\times 10^{-12}$. We study the potential observational consequences from the smallness of such a value using exact and numerical expressions for the deflection angle, optical images from optically and geometrically thin accretion disks, isoradials, and instability scales (Lyapunov index) of nearly bound geodesics associated to photon rings. Our results point towards the impossibility to distinguish between these two geometries using current and foreseeable techniques in the field of interferometric detection of optical sources.
@@ I Introduction
Despite the undeniable success of Einstein’s General Relativity (GR) in explaining a large number of gravitational phenomena in different regimes [1], there are several drawbacks of the theory. On the one hand, the accelerated expansion of the Universe, as found by Type-Ia supernovae observations  [2, 3], can only be accounted for, in the frame of GR, via a hypothetical fluid known as dark energy [4], whose properties can be assimilated to that of a cosmological constant term [5] yet there is no direct proof of its existence. Similarly, the explanation of galactic rotation curves demands more than 90% of the total mass of most galaxies to be dominated by a non-baryonic form of matter dubbed as dark matter [7], which we have been unable to detect in any terrestrial observatory yet. On more theoretical grounds, the problems with the unavoidable development of space-time singularities inside black holes and in the early cosmological evolution and the loss of predictability they bring about [6], the Hawking process and the associated information loss paradox [8], and the development of a framework unifying gravity and quantum mechanics (quantum gravity [9]), are challenges that still haunt the gravitational community.
@@
In response to these challenges, a large avenue of research has emerged in the last few decades arranged around alternative approaches to GR, generally dubbed as modified theories of gravity (see e.g. [11, 10] for some reviews). Within this context, recently Harada presented a theory called “Cotton Gravity” (CG) [12]. In this proposal, the effects describing gravity are attributed to the Cotton tensor, and the corresponding field equations contain third-order derivatives. The main interest in this proposal comes from the fact that any solution of the Einstein equations, with or without a cosmological constant term, also satisfies the field equations of CG. Furthermore, the cosmological constant itself emerges as an integration constant out of the field equations. In addition, in [13] Harada used the post-Newtonian limit of CG to solve the field equations numerically, interpreting the rotation curves of several galaxies without employing dark matter sources. Note that the possibility that dark matter may be geometric effects of modified theories of gravity has been explored in the literature [14, 15, 16, 17, 18, 19].
@@
Regarding new solutions within CG, the first non-trivial one with static, spherical symmetry was found in [13], which extends the Schwarzschild solution via a new parameter $\gamma$. Other non-trivial solutions were found in [20, 21]. On the other hand, in [22] it was emphasized that the Codazzi tensor can be derived from the equations of CG, which implies that Harada’s formulation can be understood as a modification of the GR equations due to the presence of this residual tensor. This way, with the introduction of the Codazzi tensor, it became possible to express CG through a new formulation, which allowed for the representation of conformally flat cosmological solutions without a vacuum in CG [23]. However, there is ongoing debate about the viability of CG as a potential extension of GR. Some researchers see CG as promising for addressing GR’s limitations, while others doubt its theoretical consistency [24, 25, 26, 27]. Further research is needed to determine whether CG can replace or complement GR, or if it will remain a theoretical curiosity.
@@
This paper is outlined in the following manner: In Sec. II, we briefly present the field equations of CG and the static, spherically symmetric metric used throughout this work. In Sec. III, we obtain the null geodesic equation and the effective potential of the CG black hole, using the above-mentioned spherically symmetric metric. Section IV addresses gravitational lensing in CG, where we first derive the exact analytical expression for the deflection angle using elliptic integrals, followed by an approximate expression for the weak field regime. In Sec. V, we generate the images of the Cotton black hole using the restricted $\gamma$-space parameter obtained in the previous sections, and qualitatively compare its appearance to that of a Schwarzschild black hole of the same mass parameter. In Sec. VI, we perform an analysis of isoradial curves in the Cotton and Schwarzschild space-times, to complement the black hole images presented in the previous section. In Sec. VII, we explore the Lyapunov exponent of nearly bound orbits, which is a quantity that measures the instability scale of these orbits. Finally, in Sec. VIII, we summarize our results and conclude.
@@ II A spherically symmetric solution of Cotton Gravity
Recently, as mentioned above, a generalization of the gravitational field equations of GR, dubbed as “Cotton Gravity” (CG) [12], was proposed. The field equations read as follows
@@
$-\frac{1}{6}\left(g_{\beta\mu}g_{\alpha\nu}-g_{\alpha\mu}g_{\beta\nu}\right)\Theta\,,$
with $\Theta_{\beta\mu}=g^{\alpha\nu}\Theta_{\beta\alpha\mu\nu}$ and $\Theta=g^{\beta\mu}\Theta_{\beta\mu}$.
@@
For the sake of this work we use the exact vacuum solution obtained by Harada in [13]. Similarly to GR, where $R_{\mu\nu}=0$, this solution is obtained in Cotton Gravity when $C_{\alpha\mu\nu}=0$ is considered. Such a solution belongs to the static, spherically symmetric class of metrics with line element
@@
whose metric function reads as [12]
$A(r)=1-\frac{2M}{r}-\frac{\Lambda}{3}r^{2}+\gamma r\;,\qquad\quad C(r)=r^{2}\,.$ (6)
This is the first non-trivial solution found in CG theory, which contains a new parameter $\gamma$, not present in the Schwarzschild case of the GR, and which breaks the asymptotically flat character of the space-time. Moreover, the cosmological constant is obtained naturally as an integration constant no matter the value of the cosmological constant $\Lambda$, which we take to be vanishing in this work. We draw attention to the fact that solutions containing a linear $\gamma r$ term have been obtained in other theoretical contexts (e.g. [45, 48, 49, 46, 47]).
@@ III Geodesics and effective potential
III.1 Gravitational lensing
In this section, we obtain the null geodesic equation and the effective potential of the CG black hole, using the spherically symmetric line element (5). To this end, we consider the point-particle Lagrangian density given by
@@ III.1 Gravitational lensing
where $\dot{x}^{\mu}$ is the four-velocity of the photon, $\epsilon=(-1,0,1)$ indicates the time-like, null, or space-like character of the geodesic, respectively, and an overdot denotes the derivative with respect to the affine parameter $\lambda$. Taking into account the metric (5) yields the following form for the Lagrangian
${\cal L}=\frac{1}{2}\left[-A\left(r\right)\dot{t}^{2}+\frac{\dot{r}^{2}}{A(r)}+C(r)\dot{\theta}^{2}+C(r)\sin^{2}\theta\dot{\phi}^{2}\right]=\epsilon\,.$ (8)
Given the spherically symmetric character of the space-time, we can consider the motion in the equatorial plane, $\theta=\pi/2$, without any loss of generality. In addition, since the Lagrangian (7) does not directly depend on $t$ and $\phi$, the Euler-Lagrange equations imply the conservation of two quantities:
$E$ $=A\left(r\right)\dot{t},$ (9a)
$L$ $=C\left(r\right)\dot{\phi}.,$ (9b)
where $E$ is the energy of the particle and $L$ the angular momentum, respectively (per unit mass). Since we are interested in null geodesics, $\epsilon=0$, Eq. (7) with the conserved quantities (9a) and (9b), leads to
$\dot{r}^{2}=\frac{1}{b^{2}}-V_{\rm eff}(r).$ (10)
@@
$V_{\rm eff}(r)=\frac{A(r)}{C(r)}.$ (11)
which for the metric function considered in this work, Eq. (6), reads
$V_{\rm eff}(r)=\frac{1}{r^{2}}-\frac{2M}{r^{3}}+\frac{\gamma}{r}.$ (12)
The maximum of the effective potential determines the region of unstable orbits, dubbed as the photon sphere, whose radius is denoted by $r=r_{p}$, and the corresponding impact parameter as the critical impact parameter, dubbed by $b_{c}$. The latter separates those orbits that are captured by the black hole, $b<b_{c}$, from those that are scattered back to asymptotic infinity, $b>b_{c}$. Those that have $b=b_{c}$ may linger indefinitely at the photon sphere, turning several times around the black hole. To determine these quantities, we note that circular light orbits obey $\dot{r}=0$ and $\ddot{r}=0$ simultaneously. Under these conditions, $r_{p}$ and $b_{c}$ are expressed for the metric function (6) as
$r_{p}$ $=\frac{\sqrt{1+6\gamma M}-1}{\gamma},$ (13a)
$b_{c}$ $=\frac{3\sqrt{6}\,M}{\sqrt{\left(1+\sqrt{1+6\gamma M}+6\gamma M\sqrt{1+6\gamma M}+9\gamma M\right)}},$ (13b)
which for $\gamma=0$ it recovers the critical $b_{c}=3\sqrt{3}M$ and the photon sphere radius $r_{p}=3M$ of the Schwarzschild geometry.
With the expressions (9b) and (10) we can develop the equation of motion which describes the trajectory of the photon as
$\left(\frac{dr}{d\phi}\right)^{2}=C^{2}(r)\left(\frac{1}{b^{2}}-V_{\rm eff}\right).$ (14)
These are the general expressions for a given spherically symmetric metric.
III.2 Shadow’s radius and constraints to the $\gamma$ parameter
The shadow of a black hole is a silhouette, a region marked by a strong decrease in the observed luminosity, projected onto the celestial plane of a distant observer. This projection arises from photons originating from a background source, which are absorbed due to its intense gravity. To calculate the angular size and the shadow of the black hole, we adopt the formalism described by [50]. In this approach, they consider a static observer located at ($r_{0}$, $\theta_{0}=\pi/2$), projecting an angle $\theta$ relative to the radial coordinate $r$, as given by
$\tan\theta=\sqrt{\frac{C(r)}{A^{-1}(r)}}\frac{d\phi}{dr}\Bigg{|}_{r=r_{0}}.$ (15)
@@ III.2 Shadow’s radius and constraints to the $\gamma$ parameter
To proceed with our derivation, we conveniently re-write Eq.(14) into
$\left(\frac{dr}{d\phi}\right)^{2}=\frac{C(r)}{A^{-1}(r)}\left(\frac{C(r)}{A(r)}\frac{E^{2}}{L^{2}}-1\right).$ (16)
A light ray with impact parameter $b>b_{c}$ will be scattered off the black hole at a radial distance $r_{0}>r_{m}$, while the photon sphere (i.e. the locus of unstable bound geodesics, $r_{p}$) can be found via the condition $dr/d\phi\mid_{r_{p}}=0$ which allows to express the impact parameter $b$ as
$\frac{1}{b^{2}}=\frac{A(r_{p})}{C(r_{p})}.$ (17)
Now, by introducing the definition
@@
we can rewrite Eq. (16) as
$\left(\frac{dr}{d\phi}\right)^{2}=\frac{C(r)}{A^{-1}(r)}\left(\frac{h^{2}(r)}{h^{2}(r_{p})}-1\right).$ (19)
Substituting Eq.(19) into Eq.(15), and after proper manipulations, allows to find the expression
$\sin^{2}\theta=\frac{h^{2}(r_{p})}{h^{2}(r_{0})},$ (20)
@@
Using the expressions worked out above, for CG the angular shadow’s size reads as
$\sin^{2}\theta_{s}$ $=$ $\frac{2}{r_{0}^{2}\gamma^{2}\left(\gamma+8\gamma^{2}M\right)}\Big{(}-1+6M\sqrt{\left(1+6\gamma M\right)}$ (22)
$+$ $\sqrt{\left(1+6\gamma M\right)}-9\gamma M\Big{)}\left(1-\frac{2M}{r_{0}}+\gamma r_{0}\right).$
Similarly, we can obtain an expression for the shadow radius as seen by a far-away observer, a regime in which we approximate $\sin(\theta)\sim\theta$ and $\theta_{s}=b_{c}/r_{0}$ which gives us the following relation (assuming $\theta=\pi/2$)
@@
For CG this reads as
$r_{s}=\frac{\left(\sqrt{a}-1\right)}{\gamma}\sqrt{\frac{\left(\sqrt{a}-1\right)\Bigl{[}r_{0}\left(\gamma r_{0}+1\right)-2M\Bigr{]}}{r_{0}\left(1-\sqrt{a}+4\gamma M\right)}}\,,$ (24)
where $a=1+6\gamma M$. In Fig. 1 we depict the behaviour of the shadow’s radius in CG with the parameter $\gamma$, considering the black hole distance and mass estimates reported in Table 1.
Figure 1: Shadow of a black hole in the plane in the celestial plane, expressing the celestial coordinates $(\xi,\chi)$ in a logarithmic scale, for the Schwarzschild geometry (black) and for CG with $\gamma=10^{-25},10^{-20},10^{-15}$ and $10^{-10}$ (colored and dashed lines).
We now aim to restrict the parameter space of $\gamma$ by assessing the compatibility of the CG geometry’s corresponding shadow radius $r_{s}$ with the EHT’s inferred shadow size of Sgr A* using the data above. This is so because, as argued in [52], it is reported that the observed shadow size of Sgr A* can be inferred by utilizing the size of the bright emission ring as its proxy, subject to a calibration factor that multiplicatively accounts for the theoretical and observational uncertainties. Such a correlation becomes possible thanks to priors on Sgr A*’s mass-to-distance ratio obtained via orbital dynamics measurements of the stars closest to the galactic center (i.e. dubbed S-stars), and which are reported in Table 1. These measurements allow the EHT collaboration to estimate the fractional deviation $\delta$ between Sgr A*’s inferred shadow size $r_{sh}$ and that of a Schwarzschild black hole $r_{sh,Sch}=3\sqrt{3}M$, with angular radius $\omega_{g}=M/D$ via the formula

### v2

@@ Abstract
We consider a recently introduced extension of General Relativity based on the use of the Cotton tensor and dubbed as Cotton gravity, to estimate the size of a new constant $\gamma$ appearing within a spherically symmetric, vacuum solution of such theory. Taking into account its non-asymptotically flat character, we use the inferred size of the central brightness depression of the supermassive object at the heart of the Milky Way galaxy (Sgr A*) by the Event Horizon Telescope to constrain at $2\sigma$) the CG parameter as $\gamma M\approx 3.5\times 10^{-12}$. We study the potential observational consequences from the smallness of such a value using exact and numerical expressions for the deflection angle, optical images from optically and geometrically thin accretion disks, isoradials, and instability scales (Lyapunov index) of nearly bound geodesics associated to photon rings. Our results point towards the impossibility to distinguish between these two geometries using current and foreseeable techniques in the field of interferometric detection of optical sources.
@@ I Introduction
Despite the undeniable success of Einstein’s General Relativity (GR) in explaining a large number of gravitational phenomena in different regimes [1], there are several drawbacks of the theory. On the one hand, the accelerated expansion of the Universe, as found by Type-Ia supernovae observations  [2, 3], can only be accounted for, in the frame of GR, via an hypothetical fluid known as dark energy, whose properties can be assimilated to that of a cosmological constant term yet there is no direct proof of its existence. Similarly, the explanation of galactic rotation curves demands more than 90% of the total mass of most galaxies to be dominated by a non-baryonic form of matter dubbed as dark matter, which we have been unable to detect in any terrestial observatory yet. On more theoretical grounds, the problems with the unavoidable development of space-time singularities inside black holes and in the early cosmological evolution and the loss of predictability they bring about [4], the Hawking process and the associated information loss paradox [5], and the development of a framework unifying gravity and quantum mechanics (quantum gravity [6]), are challenges that still haunt the gravitational community.
@@
In response to these challenges, a large avenue of research has emerged in the last few decades arranged around alternative approaches to GR, generally dubbed as modified theories of gravity (see e.g. [8, 7] for some reviews). Within this context, recently Harada presented a theory called “Cotton Gravity” (CG) [9]. In this proposal, the effects describing gravity are attributed to the Cotton tensor, and the corresponding field equations contain third-order derivatives. The main interest in this proposal comes from the fact that any solution of the Einstein equations, with or without a cosmological constant term, also satisfies the field equations of CG. Furthermore, the cosmological constant itself emerges as an integration constant out of the field equations. In addition, in [10] Harada used the post-Newtonian limit of CG to solve the field equations numerically, interpreting the rotation curves of several galaxies without the employ of dark matter sources.
@@
Regarding new solutions within CG, the first non-trivial one with static, spherical symmetry was found in [10], which extends the Schwarzschild solution via a new parameter $\gamma$. Such a parameter is incorporated in the vacuum solutions of the theory via the integration of its field equations, and it is interpreted as a measure of the relative strength of the contribution of the Cotton tensor to the space-time geometry. Other non-trivial solutions were found in [11, 12]. On the other hand, in [13] it was emphasized that the Codazzi tensor can be derived from the equations of CG, which implies that Harada’s formulation can be understood as a modification of the GR equations due to the presence of this residual tensor. This way, with the introduction of the Codazzi tensor, it became possible to express CG through a new formulation, which allowed for the representation of conformally flat cosmological solutions without a vacuum in CG [14].
@@ II A spherically symmetric solution of Cotton Gravity
Recently, a generalization of the field equations of GR, dubbed as “Cotton Gravity” (CG) [9], was proposed. The equations read as follows
@@
-- $\frac{1}{6}\left(g_{\beta\mu}g_{\alpha\nu}-g_{\alpha\mu}g_{\beta\nu}\right)\Theta\,,$
being $g^{\alpha\nu}\Theta_{\beta\alpha\mu\nu}=\Theta_{\beta\mu}$ and $\Theta=g^{\beta\mu}\Theta_{\beta\mu}$.
@@
For the sake of this work we shall use the exact vacuum solution obtained by Harada in [10]. Similarly to GR, where $R_{\mu\nu}=0$, this solution is obtained in Cotton Gravity when $C_{\alpha\mu\nu}=0$ is considered. Such a solution belongs to the static, spherically symmetric class of metrics with line element
@@
whose metric function, as found in [9] reads as
$A(r)=1-\frac{2M}{r}-\frac{\Lambda}{3}r^{2}+\gamma r\quad;\quad C(r)=r^{2}$ (6)
This is the first non-trivial solution found in CG theory, which contains a new parameter $\gamma$, not present in the Schwarzschild case of the GR. As long as its value is non-zero (since for $\gamma=0$ it obviously retrieves the Schwarzschld solution), this means that, even in the case of vanishing cosmological constant, $\Lambda=0$, the space-time is not asymptotically flat [^1] Such solutions, containing a term that makes the space-time neither to be asymptotically flat nor (anti-)de Sitter are not unknown in the literature, see e.g. [32, 35, 36, 33, 34]).. Given the fact that throughout the paper we shall actually assume $\Lambda=0$, then the presence of the $\gamma r$ term not only drives the asymptotic character of the geometry, but will also modify the horizons, which are now located as
$r=\frac{1\pm\sqrt{1-8M\gamma}}{2\gamma}$ (7)
where $\pm$ for the outer/inner horizons. This is more akin to the global structure of a Reissner-Nordström solution rather than a Schwarzschild one. For the purposes of this work we are solely interested in the outer (event) horizon, bearing also in mind that light propagation and deflection will be also modified in this theory. Indeed, by comparing the scales at which the gravitational mass and the $\gamma$-parameter become comparable we find a radius
$r_{h}^{\pm}\sim\sqrt{\frac{2M}{\gamma}}$ (8)
for the which $\gamma$-related effects should become relevant.
@@ III GEODESICS AND EFFECTIVE POTENTIAL
III.1 Photon’s motion in Cotton’s gravity solution
In this section, we obtain the null geodesic equation and the effective potential of the black hole described by CG, using the spherically symmetric line element and the specific solution just described above. To this end, we start our analysis by considering the point-particle Lagrangian density given by
@@ III.1 Photon’s motion in Cotton’s gravity solution
where $\dot{x}^{\mu}$ is the four-velocity of the particle, $\epsilon=(-1,0,1)$ indicates the time-like, null, or space-like character of the geodesic, respectively, and a dot denotes the derivative with respect to the affine parameter $\lambda$. The metric (5) makes the Lagrangian take the form
${\cal L}=\frac{1}{2}\left[-A\left(r\right)\dot{t}^{2}+\frac{\dot{r}^{2}}{A(r)}+C(r)\dot{\theta}^{2}+C(r)\sin^{2}\theta\dot{\phi}^{2}\right]=\epsilon,$ (10)
Frow now on we restrict our analysis to null geodesics, $\epsilon=0$. Furthermore, given the spherically symmetric character of the space-time, we can consider the motion in the equatorial plane, $\theta=\pi/2$, without any loss of generality. In addition, since the Lagrangian (9) does not directly depend on $t$ and $\phi$, the Euler-Lagrange equations imply the conservation of two quantities, $E=-A\left(r\right)\dot{t},L=C\left(r\right)\dot{\phi},$ which are the photon’s energy and angular momentum, respectively. This way, Eq.(9) with the conserved quantities above leads to
$\dot{r}^{2}=\frac{1}{b^{2}}-V_{eff}(r).$ (11)
@@
$V_{eff}(r)=\frac{A(r)}{C(r)}=\frac{1}{r^{2}}-\frac{2M}{r^{3}}+\frac{\gamma}{r}$ (12)
where the last equality correspond to the metric function considered in this work, Eq.(6). The maximum of the effective potential determines the region of unstable orbits, dubbed as the photon sphere, whose radius is denoted by $r=r_{p}$, and the corresponding impact parameter as the critical impact parameter, dubbed by $b_{c}$ (see e.g. Fig. 1).
Figure 1: Example of an unstable circular geodesic (orange) in the equatorial plane of a static and spherically symmetric black hole (black circle). In the celestial plane (blue), the critical curve (dashed red) is comprised of light rays which asymptotically approach the photon sphere.
The latter separates those orbits that are captured by the black hole, $b<b_{c}$, from those that are scattered back to asymptotic infinity, $b>b_{c}$. Those that have $b=b_{c}$ may linger indefinitely at the photon sphere, turning several times around the black hole. To determine these quantities, we note that circular light orbits obey $\dot{r}=0$ and $\ddot{r}=0$ simultaneously. Under these conditions, $r_{p}$ and $b_{c}$ are expressed for the metric function (6) as
$r_{p}$ $=\frac{\sqrt{1+6\gamma M}-1}{\gamma};$ (13a)
$b_{c}$ $=\sqrt{\frac{C(r_{p})}{r_{p}}}=\frac{3\sqrt{6}}{\sqrt{\frac{\left(1+\sqrt{1+6\gamma M}+6\gamma M\sqrt{1+6\gamma M}+9\gamma M\right)}{M^{2}}}}.$ (13b)
which recovers the critical impact parameter $b_{c}=3\sqrt{3}M$ and the photon sphere radius $r_{p}=3M$ of the Schwarzschild geometry, in the $\gamma\rightarrow 0$ limit. The geodesic equation (12) can be more conveniently written, with the help of the conserved quantities, as
$\left(\frac{dr}{d\phi}\right)^{2}=C(r)\left[\frac{1}{b^{2}}-V_{eff}\right].$ (14)
Photons whose impact parameter lie within the region $b<b_{c}$ eventually meet the event horizon. This region, bounded by $b=b_{c}$ (dubbed as the critical curve and corresponding to the projection, on the observer’s screen, of the photon sphere), is marked by a strong decrease in the observed luminosity, as projected onto the celestial plane of a distant observer. This is conventionally known in the literature as the shadow. To calculate its angular size we shall adopt the formalism described by [37]. In their approach, they assume a static observer located at ($r_{0}$, $\theta_{0}=\pi/2$) and projecting an angle $\theta$ with respect to the radial coordinate $r$ given by
$\tan\theta=\sqrt{\frac{C(r)}{A(r)}}\frac{d\phi}{dr}\Bigg{|}_{r=r_{0}}.$ (15)
@@
By introducing the definition
@@
in terms of which Eq.(14) can be written as
$\left(\frac{dr}{d\phi}\right)^{2}=\frac{C(r)}{A(r)}\left(\frac{h^{2}(r)}{h^{2}(r_{p})}-1\right).$ (17)
then by substituting into into Eq.(15) we find the expression
$sin^{2}\theta=\frac{h^{2}(r_{p})}{h^{2}(r_{0})},$ (18)
@@
For the CG solution, this angular shadow’s size reads as
$\sin^{2}\theta_{s}=\frac{2}{r_{0}^{2}\left(\gamma+8\gamma^{2}M\right)}\left(-\frac{1}{\gamma}+\frac{6M\sqrt{\left(1+6\gamma M\right)}}{1}+\frac{\sqrt{\left(1+6\gamma M\right)}}{\gamma}-9M\right)\left(-\frac{2M}{r_{0}}+\gamma r_{0}+1\right)\,.$
If, instead, we are interested in the shadow’s radius as seen by a far-away observer, for asymptotically flat space-times one can approximate $\sin(\theta)\sim\theta$ and $\theta_{s}=b_{c}/r_{0}$ so the shadow’s is simply the critical impact parameter (in units of length), that is, $r_{sh}=b_{c}=r_{p}/g_{tt}(r_{p})^{1/2}$. However, in our non-asymptotically flat space-time, we need to specify the metric functions at the specific observer radius as [37]
@@
which for CG reads as
$r_{s}=\frac{\left(\sqrt{a}-1\right)}{\gamma}\sqrt{\frac{\left(\sqrt{a}-1\right)\Bigl{[}r_{0}\left(\gamma r_{0}+1\right)-2M\Bigr{]}}{r_{0}\left(1-\sqrt{a}+4\gamma M\right)}}\,.$ (21)
where $a=1+6\gamma M$. In Fig. 2 we depict the behaviour of the shadow’s radius in CG with the parameter $\gamma$, considering the mass and distance estimates corresponding to Sgr A, reported in Table 1.
Figure 2: Black hole shadow $r_{sh}$ boundary as seen in the celestial plane $(\chi,\xi)$ of an equatorial observer, for the Schwarzschild geometry (black) and for CG with $\gamma M\in\{10^{-14},10^{-12},10^{-10}\}$ (colored and dashed lines). The celestial coordinates are computed considering the logarithm of the shadow radius (i.e. $\chi=\text{Log}_{10}\left(\frac{r_{sh}}{M}\right)\cos(\phi)$; $\xi=\text{Log}_{10}\left(\frac{r_{sh}}{M}\right)\sin(\phi)$).
III.2 Constraints to the $\gamma$ parameter from shadow’s radius
We now aim to restrict the parameter space of $\gamma$ by assessing the compatibility of the CG geometry’s corresponding shadow radius $r_{s}$ with the EHT’s inferred shadow size of Sgr A* using the data above. This is so because, as argued in [39], it is reported that the observed shadow size of Sgr A* can be inferred by utilizing the size of the bright emission ring as its proxy, subject to a calibration factor that multiplicatively accounts for the theoretical and observational uncertainties. Such a correlation becomes possible thanks to priors on Sgr A*’s mass-to-distance ratio obtained via orbital dynamics measurements of the stars closest to the galactic center (i.e. dubbed S-stars), and which are reported in Table 1. These measurements allow the EHT collaboration to estimate the fractional deviation $\delta$ between Sgr A*’s inferred shadow size $r_{h}$ and that of a Schwarzschild black hole $r_{sh,Sch}=3\sqrt{3}M$, with angular radius $\omega_{g}=M/D$ via the formula